# GPU Cluster Scheduling for Network-Sensitive Deep Learning

### v1

@@ 5.2.3. Communication overhead
Dally demonstrates the lowest communication overhead, surpassing the nearest network-aware baseline, Tiresias, by a notable margin ranging from 53% to 83% (Fig. 6 (c)). This is due to Dally’s preemption policy, which prioritizes providing better-consolidated placements to jobs suffering from sub-optimal placements or network sensitivity, thereby mitigating communication overhead. Gandiva exhibits a slight performance difference in Makespan compared to Dally in the initial 2-rack experiment but experiences a substantial degradation as the number of racks increases. This discrepancy arises because the 2-rack experiment involves fewer network connections and consequently fewer network placements. However, with an increase in the number of racks, the cluster experiences a rise in network connections and leads to jobs being placed on the network, resulting in elevated communication overhead, as depicted in Fig. 6 (c). Additionally, this phenomenon results in a long tail in the number of job completions within the cluster, subsequently leading to low cluster (GPU) utilization as depicted in Fig. 9.
@@
Figure 9. Cluster GPU utilization (8 racks)
@@
Figure 10. Number of jobs remaining (8 racks)
5.2.4. Cluster utilization
@@
We explain this section through results from the 8 racks experiment but we get similar results for the other experiments as well. Fig. 10 illustrates the number of remaining jobs in the cluster, highlighting Gandiva’s notable tailing effect. Gandiva jobs, (including the ¿P95 and ¿P99) experience substantial communication overhead, resulting in diminished cluster utilization, as depicted in Fig. 9. Dally’s efficacy in minimizing communication overhead and queuing delays contributes to the highest job completion rates among all tested schemes as shown in Fig. 10 (a proxy for cluster utilization), along with the lowest JCTs.
5.2.5. Job completion time
@@
To assess the JCT, we utilize traces with both batch and Poisson job arrivals, recognizing that the average JCT can vary significantly based on the job arrival pattern. In Fig. 11(a), we present the average JCTs with batch arrivals, revealing that Dally achieves an improvement in average JCT ranging from 19% to 36% compared to Tiresias and 23% to 51% compared to Gandiva. Examination of the JCT Cumulative Distribution Functions (CDFs) in Fig. 7 indicates that Dally reduces JCT for a majority of the job sizes. Notably, in the 2 racks experiment, Dally experiences a 39% degradation against Gandiva, which performs better with fewer aggregate network placements. However, as the number of racks (and network connections) increases, Gandiva’s performance degrades rapidly due to increased communication overhead. Additional statistics for the JCT in the 8-rack experiment are presented in Table 2 showcasing Dally’s enhancements in tail JCT.
@@
To assess Dally under Poisson arrival conditions, we conduct experiments analogous to those performed with batch arrivals, but with jobs exhibiting typical Poisson arrival times. Dally demonstrates improvements over Tiresias ranging from 16% to 34% and over Gandiva by 23% to 51% as shown in Fig. 11(b). However, in the 4-rack scenario, Dally experiences degradation compared to Tiresias due to the random arrival pattern inherent in a Poisson process. A drawback worth noting is that Tiresias exhibits superior performance in identifying optimal placements for this particular arrival pattern and network size (racks), improving upon its own average JCT in the 8-rack and 16-rack experiments. This indicates that the specific arrival pattern was well suited for Tiresias in the 4-racks case. Also, recall from Sec. 4.3.1 that Dally does not emphasize JCT when prioritizing jobs for GPU consolidation. However, Fig. 8 presents histograms of JCTs (P95) with Poisson arrival times, illustrating Dally’s improvement in the JCTs of the majority of the jobs. Detailed JCT statistics for the 8-rack Poisson experiment are provided in Table 3. We find Dally outperforming Tiresias in both average and median values by 35% and 38% respectively while being within 10% of the tail. Dally also outperforms Gandiva in average and median JCT by 83% and 88% respectively and tail JCT by 76-92%.
@@
¿P95 17485823 17841606 7097646 5983042
¿P99 39214710 25696135 8868401 8395214
Table 2. JCT batch statistics in seconds (8 racks)
@@
¿P95 71362014 15059080 39298828 16469069
¿P99 208939883 15390860 39817762 16681241
Table 3. JCT Poisson statistics in seconds (8 racks)
6. CONCLUSIONS
@@
Training of DDL jobs on GPU clusters is expensive either in a public or private cloud environment. A significant part of this time is due to the communication overhead between a hierarchical organization of GPUs in a datacenter. This communication overhead and the resultant overall training time can be minimized by a prudent allocation of GPUs, where a distributed training job is consolidated on physically proximal GPUs in proportion to its network sensitivity.
@@
In this context, we propose Dally, a cost-effective DL cluster scheduling scheme which employs delay scheduling for job placement and consolidation. Dally incorporates a novel network-sensitive preemption technique to prioritize jobs with high network sensitivity, ensuring they receive more consolidated job offers. Additionally, we propose and integrate an auto-tuner into Dally to optimize delay timers according to varying GPU contention/demand for delay scheduling. To assess the performance of Dally, we introduce ArtISt-sim, a cost-effective DL simulation platform tailored for simulating large DL clusters. Through extensive trace-driven simulations, we demonstrate that Dally can reduce end-to-end makespan time by 69% compared to consolidation based policy, average JCT by up to 83%, and communication overhead by up to 98% under congested networking conditions.
Acknowledgements
This research was supported in part by NSF grants #2116962, #2122155, #2028929.

### v2

@@ VI-3 Communication latency
Dally demonstrates the lowest communication latency, surpassing the nearest network-aware baseline, Tiresias, by a notable margin ranging from 53% to 83% (Fig. 8 (b)). This is due to Dally’s preemption policy, which prioritizes providing better-consolidated placements to jobs suffering from sub-optimal placements or network sensitivity, thereby mitigating communication latency. Gandiva exhibits a slight performance difference in makespan compared to Dally in the initial 2-rack experiment, but experiences a substantial degradation as the number of racks increases. This discrepancy arises because the 2-rack experiment involves fewer network connections and consequently fewer network placements. However, with an increase in the number of racks, the cluster experiences a rise in network connections and leads to jobs being placed on the network, resulting in elevated communication latency, as depicted in Fig. 8(b). Additionally, this phenomenon results in a long tail in the number of job completions within the cluster, subsequently leading to low cluster utilization (Fig. 11).
@@
Figure 11: Cluster GPU utilization (8 racks).
@@
Figure 12: Number of jobs remaining (8 racks).
VI-4 Cluster Utilization
@@
We explain this section through results from the 8 racks experiment but we get similar results for the other experiments as well. Fig. 12 illustrates the number of remaining jobs in the cluster, highlighting Gandiva’s notable tailing effect. Gandiva jobs (including the  P95 and  P99) experience substantial communication latency, resulting in diminished cluster utilization, as demonstrated in Fig. 11. Dally’s efficacy in minimizing communication latency and queuing delays contributes to the highest job completion rates among all tested schemes, as shown in Fig. 12 (a proxy for cluster utilization), along with the lowest JCTs. Note, in real-world scenarios, users can leverage autoscalers [aws-autoscaling] and release unused GPUs (below 100% demand) back to the cloud for cost savings. For simplicity, this work excludes the complexities of autoscaling.
VI-5 Job Completion Time (JCT)
@@
To assess the JCT, we utilize traces with both batch and Poisson job arrivals, recognizing that the average JCT can vary significantly based on the job arrival pattern. In Fig. 13(a), we present the average JCTs with batch arrivals, revealing that Dally achieves an improvement in average JCT ranging from 19% to 36% compared to Tiresias and 23% to 51% compared to Gandiva. Examination of the JCT Cumulative Distribution Functions (CDFs) in Fig. 9 indicates that Dally reduces JCT for a majority of the job sizes. Notably, in the 2 racks experiment, Dally experiences a 39% degradation against Gandiva, which performs better with fewer aggregate network placements. However, as the number of racks (and network connections) increases, Gandiva’s performance degrades rapidly due to increased communication latency. Additional statistics for the JCT in the 8-rack experiment are presented in Table II, showcasing Dally’s enhancements in tail JCT.
@@
To assess Dally under Poisson arrival conditions, we conduct experiments analogous to those performed with batch arrivals, but with jobs exhibiting typical Poisson arrival times. Dally demonstrates improvements over Tiresias ranging from 16% to 34% and over Gandiva by 23% to 51% as shown in Fig. 13(b). However, in the 4-rack scenario, Dally experiences degradation compared to Tiresias due to the random arrival pattern inherent in a Poisson process. A drawback worth noting is that Tiresias exhibits superior performance in identifying optimal placements for this particular arrival pattern and network size (racks), improving upon its own average JCT in the 8-rack and 16-rack experiments. This indicates that the specific arrival pattern was well suited for Tiresias in the 4-racks case. Also, recall from Sec. IV-B1 that Dally does not emphasize JCT when prioritizing jobs for GPU consolidation. However, Fig. 10 presents histograms of JCTs (P95) with Poisson arrival times, illustrating Dally’s improvement in the JCTs of the majority of the jobs. Detailed JCT statistics for the 8-rack Poisson experiment are provided in Table III. We find Dally outperforming Tiresias in both average and median values by 35% and 38%, respectively, while being within 10% of the tail. Dally also outperforms Gandiva in average and median JCT by 83% and 88% respectively and tail JCT by 76-92%. Note that our Poisson experiments represent peak usage scenarios to evaluate our scheme under network congestion.
@@
P95 17485823 17841606 7097646 5983042
P99 39214710 25696135 8868401 8395214
TABLE II: JCT batch statistics in seconds (8 racks).
@@
P95 71362014 15059080 39298828 16469069
P99 208939883 15390860 39817762 16681241
TABLE III: JCT Poisson statistics in seconds (8 racks).
VII RELATED WORK
VII-A DL Cluster Scheduling
A plethora of DL cluster schedulers have been developed recently with varying objectives such as JCT performance, fairness, and SLO compliance. They utilize strategies like consolidation, migration, packing, resource reallocation, etc. In terms of performance, network agnostic approaches like [lucid-asplos23, ant-man] focus on packing multiple DNN models onto a single GPU, boosting parallelism and decreasing overall execution time. Our work, centered around network sensitivity to enhance performance, is orthogonal to such efforts, as our proposed techniques can be readily applied to GPUs simultaneously running multiple models. Further, works such as [pollux, optimus, dl2, smd, aonline] improve performance through resource reallocation. However, due to the high sensitivity of DDL to network latency, these schedulers can quickly degrade performance with an unfavourable GPU placement.
[horus, muri, Co-scheML, harmony] are interference-aware schedulers and for fairness, [themis-nsdi, gandiva-fair, ASTRAEA] have been proposed. [chronus, ElasticFlow] are SLO-compliant schedulers and [sc21ones, sc23easyscale] improve resource utilization through dynamically adjusting batch size and GPU workers respectively. Recently, Sia [sia-sosp23] was introduced to reduce JCT. It employs integer linear programming (ILP) to determine the optimal GPU allocation for data-parallel deep learning jobs in a two-tier network. However, Sia overlooks the heterogeneity of network links and only accounts for two types of connections – intra-machine and inter-machine. Additionally, solving an ILP for an n-tier network could be computationally intensive. Further, to lower public cloud expenses, techniques such as [cypress, paldia, faastloop] (for DNN) and [sharma2021cash, kraken] (for data parallel workloads) have been proposed.
VII-B Traditional Cluster Scheduling
Early cluster scheduling was primarily designed for high-performance computing (HPC) workloads, leveraging the message passing interface (MPI) [open-mpi] for communication among workers. However, with the rise of big data, data centers quickly shifted towards data-parallel workloads such as MapReduce [hadoop], Spark [spark], and others. This shift led to the development of numerous job schedulers, specifically tailored for data-parallel tasks (e.g., see [delay_scheduling, DRF, grandl2014multi, grandl2016altruistic, grandl2016graphene, bhattacharya2013hierarchical, mantri-osdi10, ousterhout2017monotasks] and the references therein). Following this, need for more general-purpose schedulers capable of handling CPU-bound workloads or serving as cluster-wide schedulers became evident. This resulted in significant advancements in generic cluster scheduling. The examples include, but not limited to, [Mesos, Borg, kubernetes, schwarzkopf2013omega, ousterhout2013sparrow, delgado2016eagle]. However, deep learning workloads present unique challenges compared to both data-parallel and general CPU-bound tasks, particularly due to their reliance on exclusive GPU resource scheduling, low tolerance to network latency, and being highly repetitive.
VII-C Network Scheduling
Network scheduling deals with optimizing the network flow or the traffic in the cluster by optimizing either the network stack or the communication algorithms. They are orthogonal to job schedulers and are typically run after resources are allocated to a job. The scheduling strategy is largely driven by the datacenter workload. Some consequential work for data parallel workloads include [coflow, dens, corral, baraat, hedera]. For deep learning, recent work include [cao2024crux, rajasekaran2024cassini, mahajan2023better, li2023accelerating, lao2021atp].
VIII CONCLUSIONS
@@
Training of DDL jobs on GPU clusters is expensive in both public and private cloud environments. A significant part of this time is due to the communication latency between a hierarchical organization of GPUs in the underlying datacenter. This communication latency and the resultant overall training time can be minimized by a prudent allocation of GPUs, where a distributed training job is consolidated on physically proximal GPUs in proportion to its network sensitivity.
@@
In this context, we propose Dally, a cost-effective DL cluster scheduling scheme which employs delay scheduling for job placement and consolidation. Dally incorporates a novel network-sensitive preemption technique to prioritize jobs with high network sensitivity, ensuring they receive more consolidated job offers. Additionally, we propose and integrate an auto-tuner into Dally to optimize delay timers according to varying GPU contention/demand for delay scheduling. To assess the performance of Dally, we introduce ArtISt-sim, a cost-effective DL simulation platform tailored for simulating large DL clusters. Through extensive trace-driven simulations, we demonstrate that Dally can reduce end-to-end makespan by 69% compared to a consolidation based policy, average JCT by up to 36%, and communication latency by up to 83% under congested networking conditions.